### 21 Feb 2024 v1

# Linearity of the co-moving velocity

Abstract
The co-moving velocity is a new variable in the description of immiscible two-phase flow in porous media. It is the saturation-weighted derivatives of the seepage velocities of the two immiscible fluids with respect to saturation. I show that it is linear in the derivative of the average seepage velocity with respect to the saturation. This is in accordance with recent measurements. Two parameters are needed to describe it. In terms of relative permeability theory, this linear relation leads to a differential equation relating the two relative permeabilities describing the flow. I present this equation together with a solution.
journal: Transport in Porous Media
1 Introduction
The description of immiscible and incompressible two-phase flow in porous media at scales large enough so that the porous medium may be seen as a continuum has been a challenge for a long time. In 1936 Wyckoff and Botset wyckoff1936flow proposed a generalization of the Darcy equation darcy1856fontaines by splitting it into two independent Darcy type equations, one for each fluid, based on the idea that what each fluid will experience ,is the matrix and the other fluid together as the effective porous matrix. This reduces the effective permeability for each fluid by a factor, the relative permeability. It is then assumed that the two relative permeabilities depend only on the saturation. If I assume steady-state flow and that there are no continuum-scale gradients in the saturation, the theory is complete as described. This theory has a dominating position in all practical calculations.
The assumption that the relative permeabilies only depend on the saturation is the weak part of the theory. It is correct in a window of capillary number. However, outside this window, the pressure gradient enters and the theory loses its predictive power. For example Tallakstad et al. found that the average flow rate through a model porous medium over a rather wide range of capillary numbers depends on the pressure drop across it to a power around two tallakstad2009steady ; tallakstad2009steady2 . In their experiment, they controlled the flow rates of the two immiscible fluids entering the porous medium and could then measure the pressure drop and the saturation. But how to construct two generalized relative permeabilities from this information?
Hansen et al. hansen2018relations ; feder2022physics recognized that the problem is that knowing the flow velocity of each fluid makes it possible to find the combined average flow velocity, but the opposite is not. The combined average flow velocity must be accompanied by a second independent velocity in order to calculate the two separate fluid velocities. This would be the co-moving velocity. Roy et al. roy2022co measured the co-moving velocity from relative permeability data from multiple sources and calculated it in a dynamic pore network model finding that it would produce a straight line when plotted against the derivative of the average flow velocity with respect to the saturation, necessitating only two parameters, one setting the scale and one determining the slope. In the co-moving velocity plots generated from the relative permeability data, they observed that both parameters were constants. However, when the dynamic pore network model went outside the limits of relative permeability theory and into the non-linear flow regime, linearity of the co-moving velocity was still in place, but the two parameters would now depend on the pressure gradient. Hansen et al. hansen2023statistical constructed a statistical mechanics for immiscible two-phase flow in porous media under steady-state conditions that would explain why the derivative of the average flow velocity with respect to the saturation would be the natural variable for the co-moving velocity. Pedersen and Hansen pedersen2023parameterizations then presented an interpretation of the co-moving velocity by exploring the geometry of the space of variables describing the flow.
The last paper so far addressing the co-moving velocity is Alzubaidi et al. alzubaidi2023impact who measured it for a wide range of contact angles using lattice Boltzmann computations on the reconstructed pore spaces of North Sea Sandstone and Bentheimer Sandstone. In addition, they used experimental flow data on Bentheimer sandstone zou2018experimental . Their surprising conclusion was that the parameters characterizing the co-moving velocity were insensitive to the contact angle.
The aim of the present note is to prove that the co-moving velocity is linear when plotted against the derivative of the average flow velocity with respect to the saturation.
In the next section, I review the central concepts surrounding the co-moving velocity. In Section 3 I present the derivation of the linearity. Section 4 I explore the consequences of this linearity for relative permeability theory. In particular, I present an equation combining the two relative permeabilities. Finally, I conclude in Section 5.
2 Central concepts
I assume a porous medium sample having an area $A$ transversal to the flow direction and a porosity $\pi$. The sample is uniform. I inject under steady state conditions a volumetric flow rate $Q_{w}$ of the more wetting fluid and a volumetric flow rate $Q_{n}$ of the less wetting fluid. I assume the fluids to be well mixed in the sample so that there are no sample-scale gradients in the saturation. The total average flow rate is
$Q_{p}=Q_{w}+Q_{n}\;.$ (1)
The sub-area of the transversal area $A$ that cuts through the pores is $A_{p}=\phi A$. The sub-area of the pore area $A$ that cuts through the more wetting fluid is $A_{w}=S_{w}A_{p}$ and the sub-area that cuts through the less wetting fluid is $A_{n}=S_{n}A_{p}$.
I can then define the seepage velocities. The seepage velocity of the more wetting fluid is
$v_{w}=\frac{Q_{w}}{A_{w}}\;,$ (2)
and the less wetting fluid,
$v_{w}=\frac{Q_{n}}{A_{n}}\;.$ (3)
The combined average seepage velocity of the fluids is given by
$v_{p}=\frac{Q_{p}}{A_{p}}=S_{w}v_{w}+S_{n}v_{n}\;,$ (4)
where I in the last expression have use equations (1) – (3).
The three seepage velocities are all functions of the saturation $S_{w}$ and the pressure gradient $p=|\nabla P|$, $v_{p}=v_{p}(p,S_{w})$, $v_{w}=v_{w}(p,S_{w})$, and $v_{n}=v_{n}(p,S_{w})$.
The co-moving velocity is defined as hansen2018relations
$v_{m}=S_{w}\left(\frac{\partial v_{w}}{\partial S_{w}}\right)_{p}+S_{n}\left(\frac{\partial v_{n}}{\partial S_{w}}\right)_{p}\;.$ (5)
By taking the derivative of equation (4), I find that the co-moving velocity may be written
$v_{m}=\left(\frac{\partial v_{p}}{\partial S_{w}}\right)_{p}+v_{n}-v_{w}=v_{p}^{\prime}+v_{n}-v_{n}\;,$ (6)
where I have defined $v_{p}`=(\partial v_{p}/\partial S_{w})_{p}$.
Equations (4) and (5) constitute the mapping $(v_{w},v_{n})\to(v_{p},v_{m})$. The inverse mapping, $(v_{p},v_{m})\to(v_{w},v_{n})$, is
$v_{w}=v_{p}+S_{n}(v_{p}^{\prime}-v_{m})\;,$ (7)
$v_{n}=v_{p}-S_{w}(v_{p}^{\prime}-v_{m})\;.$ (8)
Roy et al. roy2022co and Alzubaidi et al. alzubaidi2023impact observed the co-moving velocity being linear when plotted against $v_{p}^{\prime}$,
$v_{m}=a(p)+b(p)v_{p}^{\prime}\;.$ (9)
The question I answer in the the next section is why this simple linear form?
3 Derivation
There are three steps to the derivation.
Step 1: Introduction of a velocity scale. The average seepage velocity $v_{p}$ contains an additive velocity scale that is lost when constructing $v_{p}^{\prime}$ since
[…]
where the integration constant $v_{p,0})(S_{w,0})$ is the additive velocity scale.
In the same way, the co-moving velocity $v_{m}$ will contain an additive velocity scale. This is necessary from equation (6). I will call this scale $a$ and define it as
$a=v_{m}(a,0)=[v_{n}-v_{w}]_{v_{p}^{\prime}=0}\;,$ (11)
where I from now on use the notation $v_{m}=v_{m}(a,v_{p}^{\prime})$. I also refrain from writing the pressure gradient explicitly as it is not needed explicitly in the following.
Step 2: Establishing that $v_{m}$ is an Euler homogeneous function in the velocity scale. The co-moving velocity must be invariant with respect to a change of velocity scale $v_{0}$. This was e.g., used in the analysis in Roy et al. roy2022co , as this was the basis for using the relative permeability data. This means that $v_{m}$ obeys the scaling relation
$\frac{v_{m}(a,v_{p}^{\prime})}{v_{0}}=v_{m}\left(\frac{a}{v_{0}},\frac{v_{p}^{\prime}}{v_{0}}\right)\;.$ (12)
Hence, $v_{m}$ is an Euler homogeneous function of order one in the variables $a$ and $v_{p}^{\prime}$. Taking the derivative with respect to $1/v_{0}$ and setting $v_{0}=1$ gives
$v_{m}(a,v_{p}^{\prime})=\left(\frac{\partial v_{m}}{\partial a}\right)_{v_{p}^{\prime}}a+\left(\frac{\partial v_{m}}{\partial v_{p}^{\prime}}\right)_{a}v_{p}^{\prime}\;.$ (13)
From equations (11) and (13), I must have
$\left(\frac{\partial v_{m}}{\partial a}\right)_{v_{p}^{\prime}}=1\;.$ (14)
Furthermore, I must have
$\left(\frac{\partial v_{m}}{\partial v_{p}^{\prime}}\right)_{a}=b\left(\frac{v_{p}^{\prime}}{a}\right)\;,$ (15)
where $b(v_{p}^{\prime}/a)$ is a homogeneous function of degree zero. This leads me to the following expression for $v_{m}$:
$v_{m}(a,v_{p}^{\prime})=a+b\left(\frac{v_{p}^{\prime}}{a}\right)\ v_{p}^{\prime}\;.$ (16)
Step 3: Combining the two previous steps. I now combine equations (14), (15) and (16),
$\left(\frac{\partial v_{m}}{\partial a}\right)_{v_{p}^{\prime}}=\left(\frac{\partial}{\partial a}\right)_{v_{p}^{\prime}}\left[a+b\left(\frac{v_{p}^{\prime}}{a}\right)\ v_{p}^{\prime}\right]=1-\frac{db(u)}{du}\bigg{|}_{u=v_{p}^{\prime}/a}\left(\frac{v_{p}^{\prime}}{a}\right)^{2}=1\;,$ (17)
where the rightmost equality is equation (14). I have here defined the new combined variable $u=v_{p}^{\prime}/a$. I see from equation (17) that
$\frac{db(u)}{du}=0\;.$ (18)
Hence, $b$ is a constant. Combining this result with equation (16) gives the final result,
$v_{m}(a,v_{p}^{\prime})=a+b\ v_{p}^{\prime}\;.$ (19)
The co-moving velocity $v_{m}$ is linear in $v_{p}^{\prime}$.
4 Consequences for relative permeability theory
The generalized Darcy equations that forms the core of relative permeability theory are
$v_{w}=-\frac{Kk_{rw}}{\mu_{w}\phi S_{w}}\ p\;,$ (20)
$v_{n}=-\frac{Kk_{rn}}{\mu_{n}\phi S_{n}}\ p\;,$ (21)
where $\mu_{w}$ and $\mu_{n}$ are viscosities, $K$ the permeability and $k_{rw}(S_{w})$ and $k_{rn}(S_{w})$ are the relative permeabilities. The saturations appear in the expressions since $v_{w}$ and $v_{n}$ are the seepage velocities.
Following Roy et al. roy2022co , I define a velocity scale
$v_{0}=-\frac{K}{\mu_{w}\phi}\ p\;,$ (23)
which is the seepage velocity of the wetting fluid at saturation $S_{w}=1$. I may then write the generalized Darcy equations as
$v_{w}$ $=$ $\frac{k_{rw}}{S_{w}}\ v_{0}\;,$ (24)
$v_{n}$ $=$ $\frac{k_{rn}}{S_{n}}\ Mv_{0}\;,$ (25)
where I have defined the viscosity ratio $M=\mu_{w}/\mu_{n}$. I now combine these two expressions with the co-moving velocity, equation (19) using equation (6). I also make a change in notation by setting $a\to av_{0}$, as in roy2022co . The resulting equation is
$(1-b)\frac{d}{dS_{w}}\left[k_{rw}+Mk_{rn}\right]-\frac{k_{rw}}{S_{w}}+\frac{Mk_{rn}}{S_{n}}=a\;.$ (27)
A simplified version of this equation was used in Alzubaidi et al. alzubaidi2023impact to determine the non-wetting relative permeability when the wetting permeability, $a=0$ and $b$ is known. One way to use this equation would be to plug the known $k_{rw}$ into it and then find $k_{rn}$ through numerical integration.
It is also possible to find analytic solutions to equation (27). I give an example in the following. Picci and Battiato picchi2019relative suggested the semi-empirical form (note that their viscosity ratio is the inverse of the definition I use)
$k_{rw}$ $=$ $k_{rw}^{0}S_{w}^{2}\;,$ (28)
$k_{rn}$ $=$ $k_{rn}^{0}S_{n}^{2}\left[1+\frac{2S_{w}}{MS_{n}}\right]\;,$ (29)
where $k_{rw}^{0}$ and $k_{rn}^{0}$ are the end point relative permeabilities. Setting
$b=\frac{1}{2}\;,$ (30)
and
$a=k_{rn}^{0}$ (31)
solves equation (27).
By solving for the saturation for which $v_{p}^{\prime}=0$, using equations (28) and (29), we find that
$[v_{n}-v_{w}]_{v_{p}^{\prime}=0}=k_{rn}^{0}\;.$ (32)
According to equation (11) this is equal to $a$, as found solving equation (27) directly, leading to (31).
The Corey empirical relative permeabilities are, on the other hand, not solutions to equation (27). That is, when the Corey exponents are equal, they are solutions, but with an unphysical $a=0$. When the exponents are different, they are not solutions to the equation for any combination of $b$ and $a$.
5 Discussion and Conclusion
I have in this paper shown that the co-moving velocity is of the form given by equation (19). The derivation is based on two assumptions, 1. the necessity of an additive velocity scale $a$, and 2. the invariance of $v_{m}$ with respect to a change of scale $v_{0}$. Both assumptions are weak and does not hinge on any restrictions on the constitutive law for $v_{p}$, linear or non-linear. In order to connect relative permeability theory to the formalism for which $v_{m}$ is a central variable, the specific constitutive law
$v_{p}(p,S_{w})=\left[k_{rw}(S_{w})+Mk_{rn}(S_{w})\right]v_{0}(p)\;,$ (33)
is used. Equation (6), which is the combination of the relative permeability constitutive laws with the linear equation for $v_{m}$, (19), leads to a relation between the two relative permeabilities: knowing one, gives the other one.
The parameter $a=[v_{n}-v_{w}]_{v_{p}^{\prime}=0}$, equation (11) is understood. However, the second parameter, $b$ is not. It typically takes on a value around 0.6–0.8 and seems quite insensitive to the contact angle between the fluids and the matrix as shown by Alzubaidi et al. alzubaidi2023impact . It is an open question why. Furthermore, from equation (6) one should think that $b=1$. It is not. Why?
Declaration — The author declares no conflict of interest.
Acknowledgement — The author thanks R. T. Armstrong, C. F. Berg, S. Berg, E. G. Flekkøy, J. E. McClure, B. Hafskjold, H. Pedersen, T. Ramstad, S. Roy, and S. Sinha for countless discussions. This work was partly supported by the Research Council of Norway through its Centres of Excellence funding scheme, project number 262644.